# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 1. Introduction
In this paper, we initiate the study of abelian pro-Lie Polish abelian groups from the viewpoint of homological algebra. Pro-Lie groups [19] are the topological groups that can be written as limits of an inverse system of Lie groups. We will focus on pro-Lie groups that are furthermore Polish, in which case they can written as limits of an inverse sequence of Lie groups. The class of pro-Lie Polish groups is a natural extension of the class of Lie groups, and it satisfies desirable closure properties, as it is closed under countable products, closed subgroups, and quotients by closed subgroups within the class of Polish groups.
We will focus on the case of abelian pro-Lie Polish abelian groups, as they form a category enriched over abelian groups, where the machinery from homological algebra can be applied. The category of abelian Polish pro-Lie groups contains all the locally compact Polish abelian groups, as well as the non-Archimedean Polish abelian groups (which are inverse limits of countable discrete groups). In fact, the class of abelian pro-Lie Polish abelian groups is the smallest class of Polish abelian groups that contains all locally compact Polish abelian groups and non-Archimedean Polish abelian groups and it is closed under extensions.
The study of homological invariants in the context of topological abelian groups followed rapidly their introduction in the purely algebraic setting, going as far back as the 1940s [9, 30]; see also [8, 26, 6, 12, 18, 17, 29]. This study has been most commonly undertaken for locally compact topological Hausdorff (Polish) abelian groups. Injective and projective objects in this category were characterized by Moskowitz in [29], showing in particular that this category does not have enough injective and projective objects. The Yoneda $\mathrm{Ext}$ groups in terms of extensions were introduced and studied by Fulp and Griffith in [12], where it is proved that $\mathrm{Ext}^{2}$ vanishes. The functor $\mathrm{Ext}$ for locally compact Hausdorff topological abelian groups was later studied from the viewpoint of derived categories by Hoffmann and Spitweck [21], who recognized it as the cohomological right derived functor of $\mathrm{Hom}$ in the sense of Verdier [36], despite the category of locally compact Hausdorff topological groups not having enough injectives or projectives.
@@
The rest of this paper is divided into three sections. In Section 2 we recall some preliminary notions from category theory, including the notions of exact and quasi-abelian category and the corresponding notions of derived category and functor. In Section 3 we obtain our main results outlined above concerning the category pro-Lie of Polish abelian groups. Finally, in Section 4 we consider several natural thick (or just fully exact) subcategories of $\mathbf{proLiePAb}$, characterizing injective and projective objects in each of them, and determining whether they have enough injectives or projectives.
@@ 2.1. Exact categories
An additive category [28, Section VIII.2] is a category enriched over the category $\mathbf{Ab}$ of abelian groups that also has a terminal object (which is necessarily also initial, and called the zero object) and binary products (which are necessarily also coproducts, and called biproducts). A kernel-cokernel pair in an additive category $\mathcal{A}$ is a pair $\left(f,g\right)$ of arrows in $\mathcal{A}$ such that $f$ is the kernel of $g$ and $g$ is the cokernel of $f$. This can be seen as an object in the double arrow category $\mathcal{A}^{\cdot\rightarrow\cdot\rightarrow\cdot}$ of $\mathcal{A}$.
@@
In the context of finitely complete and finitely cocomplete categories, a stronger notion of exact functor is considered. If $\mathcal{C}$ and $\mathcal{D}$ are categories that are finitely complete and finitely cocomplete (i.e., have finite limits and finite colimits), then a functor $F:\mathcal{C}\rightarrow\mathcal{D}$ is called exact if it is finitely continuous and finitely cocontinuous (i.e., commutes with finite limits and finite colimits); see [24, Section 3.3]. An exact functor in the sense of [7, Proposition 5.2] between finitely complete and finitely cocomplete exact categories is not necessarily exact in the sense of [24, Section 3.3].
@@ 2.2. Derived categories
A morphism $f:A\rightarrow A$ in $\mathcal{A}$ is idempotent if $f=f\circ f$. The category $\mathcal{A}$ is idempotent-complete if every idempotent morphism has a kernel. Suppose now that $\mathcal{A}$ is an idempotent-complete exact category. A complex $A$ over $\mathcal{A}$ is acyclic [7, Section 10] if it has admissible differentials and it is exact at each degree as in Definition 2.3. We let $\mathrm{N}\left(\mathcal{A}\right)$ be the full subcategory of $\mathrm{K}\left(\mathcal{A}\right)$ spanned by acyclic complexes. Then we have that $\mathrm{N}\left(\mathcal{A}\right)$ is a thick subcategory [31, Definition 2.1.6] of the triangulated category $\mathrm{K}\left(\mathcal{A}\right)$ [7, Corollary 10.11]. This means that $\mathrm{N}\left(\mathcal{A}\right)$ is a triangulated subcategory of $\mathrm{K}\left(\mathcal{A}\right)$ [31, Definition 1.5.1] and contains all the direct summands of its objects,
One can thus define the corresponding Verdier quotient $\mathrm{D}\left(\mathcal{A}\right):=\mathrm{K}\left(\mathcal{A}\right)/\mathrm{N}\left(\mathcal{A}\right)$ [31, Theorem 2.1.8], which is called the derived category of $\mathcal{A}$ [7, Section 10]. By definition, this is the localization $\mathrm{K}\left(\mathcal{A}\right)[\Sigma^{-1}]$ where $\Sigma$ is the multiplicative system in $\mathrm{K}\left(\mathcal{A}\right)$ consisting of homotopy classes of chain maps whose mapping cone is acyclic, called quasi-isomorphisms [7, Definition 10.16]. If $Q_{\mathcal{A}}:\mathrm{K}\left(\mathcal{A}\right)\rightarrow\mathrm{D}\left(\mathcal{A}\right)$ is the quotient functor, then a morphism in $\mathrm{K}\left(\mathcal{A}\right)$ is a quasi-isomorphism if and only if its image in $\mathrm{D}\left(\mathcal{A}\right)$ is invertible, i.e. the multiplicative system $\Sigma$ is saturated [7, Remark 10.19].
@@ 2.3. Abelian and quasi-abelian categories
Let $\mathcal{A}$ be a quasi-abelian category, and $\mathcal{B}$ be a (not necessarily full) subcategory of $\mathcal{A}$. Then we say that $\mathcal{B}$ is a quasi-abelian subcategory of $\mathcal{A}$ if $\mathcal{B}$ is an abelian category and the inclusion functor $\mathcal{B}\rightarrow\mathcal{A}$ is finitely continuous and finitely cocontinuous. We say that $\mathcal{B}$ is a fully exact quasi-abelian subcategory if it is a quasi-abelian subcategory that is also a fully exact subcategory (where $\mathcal{A}$ is endowed with its canonical exact structure). In this case, we also say that $\mathcal{B}$ is a thick subcategory of $\mathcal{A}$; see [24, Definition 8.3.21]. (Notice that this is different from the notion of thick subcategory of a triangulated category from [31, Definition 2.1.6].)
@@
Let $\mathcal{A}$ be a quasi-abelian category and $\mathrm{D}\left(\mathcal{A}\right)$ be its derived category of $\mathcal{A}$. The left heart of $\mathcal{A}$ is by definition the heart of $\mathrm{D}\left(\mathcal{A}\right)$ with respect to its canonical left truncation structure [34, Definition 1.2.18]. Concretely, $\mathrm{LH}\left(\mathcal{A}\right)$ is the full subcategory spanned by complexes $A$ with $A^{n}=0$ for $n\in\mathbb{Z}\setminus\left\{-1,0\right\}$ and such that the differential $A^{-1}\rightarrow A^{0}$ is monic. Then $\mathrm{LH}\left(\mathcal{A}\right)$ is an abelian category that contains $\mathcal{A}$ as a thick subcategory [34, Proposition 1.2.29]. Furthermore, the inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ satisfies the following universal property: for every abelian category $\mathcal{M}$ and finitely continuous and exact functor $F:\mathcal{A}\rightarrow\mathcal{M}$ there exists an essentially unique functor $\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{M}$ whose restriction to $\mathcal{A}$ is isomorphic to $F$ [34, Proposition 1.2.34]. The inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ extends to an equivalence of categories $\mathrm{D}\left(\mathcal{A}\right)\rightarrow\mathrm{D}\left(\mathrm{LH}\left(\mathcal{A}\right)\right)$ [34, Proposition 1.2.32], and it has a left adjoint $\kappa:\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{A}$ [34, Definition 1.2.26 and Proposition 1.2.27].
@@
The canonical left truncation structure on $\mathrm{D}\left(\mathcal{A}\right)$ yields the cohomology functors $\mathrm{H}^{n}:\mathrm{D}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ for $n\in\mathbb{Z}$ [34, Definition 1.2.18]. We have that a morphism $f$ in $\mathrm{D}\left(\mathcal{A}\right)$ is an isomorphism if and only if $\mathrm{H}^{n}\left(f\right)$ is an isomorphism for every $n\in\mathbb{Z}$.
@@
Recall that a torsion pair [35, Definition 2.4] in a quasi-abelian category $\mathcal{M}$ is a pair $\left(\mathcal{T},\mathcal{F}\right)$ of full subcategories, where $\mathcal{T}$ is called a torsion class and $\mathcal{F}$ called a torsion-free class, such that:
@@
    for all objects $T$ of $\mathcal{T}$ and $A$ of $\mathcal{F}$, $\mathrm{Hom}\left(T,F\right)=0$;
@@
In this case, we have that an object $X$ of $\mathcal{M}$ is in $\mathcal{T}$ if and only if $\mathrm{Hom}\left(X,C\right)=0$ for all $C$ in $\mathcal{F}$, and it is in $\mathcal{F}$ if and only if $\mathrm{Hom}\left(T,X\right)=0$ for all $T$ in $\mathcal{T}$. If $\left(\mathcal{T},\mathcal{F}\right)$ is a torsion pair for $\mathcal{M}$, then $\mathcal{T}$ and $\mathcal{F}$ are full subcategories of $\mathcal{M}$ (essentially) closed under extensions [33, Theorem 2]. Conversely, if $\mathcal{A}$ is a quasi-abelian category, let $\kappa:\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{A}$ be the left adjoint of the inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$. Thus, we have that
@@
for all objects $B$ of $\mathcal{A}$. Define $\mathcal{T}$ to be the full subcategory of $\mathrm{LH}\left(\mathcal{A}\right)$ spanned by the objects $X$ such that $\kappa\left(X\right)=0$, and define $\mathcal{F}$ to be equal to $\mathcal{A}$. Then we have that $\left(\mathcal{T},\mathcal{F}\right)$ is a torsion pair in $\mathrm{LH}\left(\mathcal{A}\right)$ [33, Theorem 2].
@@
Let $\mathcal{M}$ be an abelian category. A quasi-abelian $\mathcal{M}$-enriched category $\mathcal{A}$ is a quasi-abelian category that is also $\mathcal{M}$-enriched, and such that finite limit and finite colimit functors are also $\mathcal{M}$-enriched.
If $\mathcal{C}$ is a category, we let $\mathrm{pro}\left(\mathcal{C}\right)$ be the pro-category of $\mathcal{C}$ [24, Definition 6.1.1]. We let $\mathrm{pro}_{\omega}\left(\mathcal{C}\right)$ be the full subcategory of $\mathrm{pro}\left(\mathcal{C}\right)$ spanned by functors $\beta:D^{\mathrm{op}}\rightarrow\mathcal{C}$ where $D$ is a countable directed ordered set. If $\mathcal{A}$ is a quasi-abelian category, then $\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$ is also a quasi-abelian category [24, Theorem 8.6.5]. Furthermore, $\mathcal{A}$ is a thick subcategory of $\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$ [24, Proposition 8.6.11]. The same proof as [24, Lemma 8.6.4 and Proposition 8.6.12] gives the following.
Lemma 2.4.
Suppose that $\mathcal{A}$ is a quasi-abelian category, and $\mathcal{B}$ is a full subcategory of $\mathcal{A}$. We have that:
  • $\mathcal{B}$ is closed under quotients in $\mathcal{A}$ if and only if $\mathrm{pro}_{\omega}\left(\mathcal{B}\right)$ is closed under quotients in $\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$;
  • $\mathcal{B}$ is closed under subobjects in $\mathcal{A}$ if and only if $\mathrm{pro}_{\omega}\left(\mathcal{B}\right)$ is closed under subobjects in $\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$;
  • $\mathcal{B}$ is closed under extensions in $\mathcal{A}$ if and only if $\mathrm{pro}_{\omega}\left(\mathcal{B}\right)$ is closed under extensions in $\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$.

### v2

@@ 1. Introduction
In this paper, we initiate the study of pro-Lie Polish abelian groups from the viewpoint of homological algebra. Pro-Lie groups [20] are the topological groups that can be written as limits of an inverse system of Lie groups. We will focus on pro-Lie groups that are furthermore Polish, in which case they can written as limits of an inverse sequence of Lie groups. The class of pro-Lie Polish groups is a natural extension of the class of Lie groups, and it satisfies desirable closure properties, as it is closed under countable products, closed subgroups, and quotients by closed subgroups within the class of Polish groups.
We will focus on the case of abelian pro-Lie Polish groups, as they form a category enriched over abelian groups, where the machinery from homological algebra can be applied. The category of abelian Polish pro-Lie groups contains all the locally compact Polish abelian groups, as well as the non-Archimedean Polish abelian groups (which are inverse limits of countable discrete groups). In fact, the class of pro-Lie Polish abelian groups is the smallest class of Polish abelian groups that contains all locally compact Polish abelian groups and non-Archimedean Polish abelian groups and it is closed under extensions.
The study of homological invariants in the context of topological abelian groups followed rapidly their introduction in the purely algebraic setting, going as far back as the 1940s [9, 31]; see also [8, 27, 6, 12, 18, 17, 30]. This study has been most commonly undertaken for locally compact topological Hausdorff (Polish) abelian groups. Injective and projective objects in this category were characterized by Moskowitz in [30], showing in particular that this category does not have enough injective and projective objects. The Yoneda $\mathrm{Ext}$ groups in terms of extensions were introduced and studied by Fulp and Griffith in [12], where it is proved that $\mathrm{Ext}^{2}$ vanishes. The functor $\mathrm{Ext}$ for locally compact Hausdorff topological abelian groups was later studied from the viewpoint of derived categories by Hoffmann and Spitzweck [22], who recognized it as the cohomological right derived functor of $\mathrm{Hom}$ in the sense of Verdier [37], despite the category of locally compact Hausdorff topological groups not having enough injectives or projectives.
@@
The rest of this paper is divided into three sections. In Section 2 we recall some preliminary notions from category theory, including the notions of exact and quasi-abelian category and the corresponding notions of derived category and functor. In Section 3 we obtain our main results outlined above concerning the category of pro-Lie Polish abelian groups. Finally, in Section 4 we consider several natural thick (or just fully exact) subcategories of $\mathbf{proLiePAb}$, characterizing injective and projective objects in each of them, and determining whether they have enough injectives or projectives.
Acknowledgments
We thank Jeffrey Bergfalk, André Nies, and Joe Zielinski for many useful conversations, and Luca Marchiori for a careful reading of a preliminary version of this paper and a large number of helpful suggestions.
@@ 2.1. Exact categories
An additive category [29, Section VIII.2] is a category enriched over the category $\mathbf{Ab}$ of abelian groups that also has a terminal object (which is necessarily also initial, and called the zero object) and binary products (which are necessarily also coproducts, and called biproducts). A kernel-cokernel pair in an additive category $\mathcal{A}$ is a pair $\left(f,g\right)$ of arrows in $\mathcal{A}$ such that $f$ is the kernel of $g$ and $g$ is the cokernel of $f$. This can be seen as an object in the double arrow category $\mathcal{A}^{\cdot\rightarrow\cdot\rightarrow\cdot}$ of $\mathcal{A}$.
@@
In the context of finitely complete and finitely cocomplete categories, a stronger notion of exact functor is considered. If $\mathcal{C}$ and $\mathcal{D}$ are categories that are finitely complete and finitely cocomplete (i.e., have finite limits and finite colimits), then a functor $F:\mathcal{C}\rightarrow\mathcal{D}$ is called exact if it is finitely continuous and finitely cocontinuous (i.e., commutes with finite limits and finite colimits); see [25, Section 3.3]. An exact functor in the sense of [7, Proposition 5.2] between finitely complete and finitely cocomplete exact categories is not necessarily exact in the sense of [25, Section 3.3].
@@ 2.2. Derived categories
A morphism $f:A\rightarrow A$ in $\mathcal{A}$ is idempotent if $f=f\circ f$. The category $\mathcal{A}$ is idempotent-complete if every idempotent morphism has a kernel. Suppose now that $\mathcal{A}$ is an idempotent-complete exact category. A complex $A$ over $\mathcal{A}$ is acyclic [7, Section 10] if it has admissible differentials and it is exact at each degree as in Definition 2.3. We let $\mathrm{N}\left(\mathcal{A}\right)$ be the full subcategory of $\mathrm{K}\left(\mathcal{A}\right)$ spanned by acyclic complexes. Then we have that $\mathrm{N}\left(\mathcal{A}\right)$ is a thick subcategory [32, Definition 2.1.6] of the triangulated category $\mathrm{K}\left(\mathcal{A}\right)$ [7, Corollary 10.11]. This means that $\mathrm{N}\left(\mathcal{A}\right)$ is a triangulated subcategory of $\mathrm{K}\left(\mathcal{A}\right)$ [32, Definition 1.5.1] and contains all the direct summands of its objects.
One can thus define the corresponding Verdier quotient $\mathrm{D}\left(\mathcal{A}\right):=\mathrm{K}\left(\mathcal{A}\right)/\mathrm{N}\left(\mathcal{A}\right)$ [32, Theorem 2.1.8], which is called the derived category of $\mathcal{A}$ [7, Section 10]. By definition, this is the localization $\mathrm{K}\left(\mathcal{A}\right)[\Sigma^{-1}]$ where $\Sigma$ is the multiplicative system in $\mathrm{K}\left(\mathcal{A}\right)$ consisting of homotopy classes of chain maps whose mapping cone is acyclic, called quasi-isomorphisms [7, Definition 10.16]. If $Q_{\mathcal{A}}:\mathrm{K}\left(\mathcal{A}\right)\rightarrow\mathrm{D}\left(\mathcal{A}\right)$ is the quotient functor, then a morphism in $\mathrm{K}\left(\mathcal{A}\right)$ is a quasi-isomorphism if and only if its image in $\mathrm{D}\left(\mathcal{A}\right)$ is invertible, i.e. the multiplicative system $\Sigma$ is saturated [7, Remark 10.19].
@@ 2.3. Abelian and quasi-abelian categories
Let $\mathcal{A}$ be a quasi-abelian category, and $\mathcal{B}$ be a (not necessarily full) subcategory of $\mathcal{A}$. Then we say that $\mathcal{B}$ is a quasi-abelian subcategory of $\mathcal{A}$ if $\mathcal{B}$ is a quasi-abelian category and the inclusion functor $\mathcal{B}\rightarrow\mathcal{A}$ is finitely continuous and finitely cocontinuous. We say that $\mathcal{B}$ is a fully exact quasi-abelian subcategory if it is a quasi-abelian subcategory that is also a fully exact subcategory (where $\mathcal{A}$ is endowed with its canonical exact structure). In this case, we also say that $\mathcal{B}$ is a thick subcategory of $\mathcal{A}$; see [25, Definition 8.3.21]. (Notice that this is different from the notion of thick subcategory of a triangulated category from [32, Definition 2.1.6].)
@@
Let $\mathcal{A}$ be a quasi-abelian category and $\mathrm{D}\left(\mathcal{A}\right)$ be its derived category. The left heart of $\mathcal{A}$ is by definition the heart of $\mathrm{D}\left(\mathcal{A}\right)$ with respect to its canonical left truncation structure [35, Definition 1.2.18]. Concretely, $\mathrm{LH}\left(\mathcal{A}\right)$ is the full subcategory spanned by complexes $A$ with $A^{n}=0$ for $n\in\mathbb{Z}\setminus\left\{-1,0\right\}$ and such that the differential $A^{-1}\rightarrow A^{0}$ is monic. Then $\mathrm{LH}\left(\mathcal{A}\right)$ is an abelian category that contains $\mathcal{A}$ as a thick subcategory [35, Proposition 1.2.29]. Furthermore, the inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ satisfies the following universal property: for every abelian category $\mathcal{M}$ and finitely continuous and exact functor $F:\mathcal{A}\rightarrow\mathcal{M}$ there exists an essentially unique functor $\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{M}$ whose restriction to $\mathcal{A}$ is isomorphic to $F$ [35, Proposition 1.2.34]. The inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ extends to an equivalence of categories $\mathrm{D}\left(\mathcal{A}\right)\rightarrow\mathrm{D}\left(\mathrm{LH}\left(\mathcal{A}\right)\right)$ [35, Proposition 1.2.32], and it has a left adjoint $\kappa:\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{A}$ [35, Definition 1.2.26 and Proposition 1.2.27].
@@
The canonical left truncation structure on $\mathrm{D}\left(\mathcal{A}\right)$ yields the cohomology functors $\mathrm{H}^{n}:\mathrm{D}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$ for $n\in\mathbb{Z}$ [35, Definition 1.2.18]. We have that a morphism $f$ in $\mathrm{D}\left(\mathcal{A}\right)$ is an isomorphism if and only if $\mathrm{H}^{n}\left(f\right)$ is an isomorphism for every $n\in\mathbb{Z}$.
@@
Recall that a torsion pair [36, Definition 2.4] in a quasi-abelian category $\mathcal{M}$ is a pair $\left(\mathcal{T},\mathcal{F}\right)$ of full subcategories, where $\mathcal{T}$ is called a torsion class and $\mathcal{F}$ called a torsion-free class, such that:
@@
    for all objects $T$ of $\mathcal{T}$ and $F$ of $\mathcal{F}$, $\mathrm{Hom}\left(T,F\right)=0$;
@@
In this case, we have that an object $X$ of $\mathcal{M}$ is in $\mathcal{T}$ if and only if $\mathrm{Hom}\left(X,C\right)=0$ for all $C$ in $\mathcal{F}$, and it is in $\mathcal{F}$ if and only if $\mathrm{Hom}\left(T,X\right)=0$ for all $T$ in $\mathcal{T}$. If $\left(\mathcal{T},\mathcal{F}\right)$ is a torsion pair for $\mathcal{M}$, then $\mathcal{T}$ and $\mathcal{F}$ are full subcategories of $\mathcal{M}$ (essentially) closed under extensions [34, Theorem 2]. Conversely, if $\mathcal{A}$ is a quasi-abelian category, let $\kappa:\mathrm{LH}\left(\mathcal{A}\right)\rightarrow\mathcal{A}$ be the left adjoint of the inclusion $\mathcal{A}\rightarrow\mathrm{LH}\left(\mathcal{A}\right)$. Thus, we have that
@@
for all objects $B$ of $\mathcal{A}$. Define $\mathcal{T}$ to be the full subcategory of $\mathrm{LH}\left(\mathcal{A}\right)$ spanned by the objects $X$ such that $\kappa\left(X\right)=0$, and define $\mathcal{F}$ to be equal to $\mathcal{A}$. Then we have that $\left(\mathcal{T},\mathcal{F}\right)$ is a torsion pair in $\mathrm{LH}\left(\mathcal{A}\right)$ [34, Theorem 2].
Recall that if $\mathcal{C}$ is a category with finite products and initial object $1$, then an (abelian) group in a category is a tuple $\left(G,m,i,u\right)$ where $G$ is an object of $\mathcal{C}$, $m:G\times G\rightarrow G$, $i:G\rightarrow G$, and $u:1\rightarrow G$ are morphisms in $\mathcal{C}$ that satisfy the “group axioms” (and the commutativity of the operation), expressed in terms of morphisms of $\mathcal{C}$ [2, Definition 4.1]. A homomorphism between group objects in $\mathcal{C}$ is a morphism in $\mathcal{C}$ between group objects that commutes with the “group operations” [2, Definition 4.2]. In the category of (abelian) groups, every object has a unique group structure, which is automatically abelian [2, Proposition 4.5].
Definition 2.4.
A category of abelian groups is a category $\mathcal{C}$ with finite products such that every object of $\mathcal{C}$ has a unique abelian group structure.
We regard an object of a category of abelian groups as an abelian group object in $\mathcal{C}$ with respect to its unique abelian group structure. Notice that a morphism in $\mathcal{C}$ is automatically a homomorphism between group objects, with respect to such a unique abelian group structure.
For example, the category of abelian Polish groups is a category of abelian groups. Likewise, any quasi-abelian subcategory of the category of abelian Polish groups is a category of abelian groups. If $R$ is a ring, then the category of (left) $R$-modules is a category of abelian groups.
Recall that an $\mathbf{Ab}$-category is a category enriched over the (monoidal) category $\mathbf{Ab}$ of abelian groups. An additive category is an $\mathbf{Ab}$-category with that has a zero object and biproducts [29, Chapter VIII].
Suppose that $\mathcal{M}$ is a category of abelian groups. For objects $A,B$ of $\mathcal{M}$ and morphisms $\varphi,\psi:A\rightarrow B$ define $\varphi+\psi:A\rightarrow B$ to be the morphism $m_{B}\circ\left(\varphi\times\psi\right)\circ\Delta_{A}$ where $\Delta_{A}:A\rightarrow A\times A$ is the canonical diagonal morphism, $m_{B}$ is the group operation in $B$.
Proposition 2.5.
Let $\mathcal{M}$ be a locally small category of abelian groups. Then $\mathcal{M}$ is an additive category with respect to the operation on $\mathrm{Hom}$-sets defined above.
Proof.
Associativity of the operation follows from associativity of $m_{B}$. It is easily seen that the unique morphism $0:A\rightarrow B$ that factors through the initial object of $\mathcal{M}$ is the neutral element with respect to the operation, while $i_{B}\circ\varphi$ is the opposite of a morphism $\varphi$. Since by assumption $\mathcal{M}$ has all finite products, it is an additive category [29, Chapter VIII, Section 2, Proposition 1 and Theorem 2]. ∎
We define a (quasi-)abelian category of abelian groups to be a locally small category of abelian groups that is (quasi-)abelian with respect to the additive category structure as in the previous proposition.
We would like to consider the natural generalization of $\mathbf{Ab}$-category where $\mathbf{Ab}$ is replaced with a category of abelian groups that is not necessarily monoidal, such as the category of pro-Lie Polish abelian groups.
Definition 2.6.
Let $\mathcal{M}$ be a quasi-abelian category of abelian groups. An $\mathcal{M}$-category is a category $\mathcal{A}$ such that for every object $A$ and $B$ of $\mathcal{A}$, $\mathrm{Hom}_{\mathcal{A}}\left(A,B\right)$ is an object of $\mathcal{M}$, and $\mathrm{Hom}_{\mathcal{A}}\left(-,B\right)$ an $\mathrm{Hom}_{\mathcal{A}}\left(A,-\right)$ are functors $\mathcal{A}\rightarrow\mathcal{M}$. Given $\mathcal{M}$-categories $\mathcal{A}$ and $\mathcal{B}$, an additive $\mathcal{M}$-functor $F:\mathcal{A}\rightarrow\mathcal{B}$ is a functor such that, for every object $A,B$ in $\mathcal{A}$, $\mathrm{Hom}\left(A,B\right)\rightarrow\mathrm{Hom}\left(FA,FB\right)$ is a morphism in $\mathcal{M}$.
Example 2.7.
If $R$ is a commutative ring, then $R$-modules form an abelian category of abelian groups $\mathrm{Mod}\left(R\right)$. A $\mathrm{Mod}\left(R\right)$-category in the sense of Definition 2.6 is the same as an $R$-linear category.
Notice that, if $\mathcal{A}$ is an $\mathcal{M}$-category, then $\mathcal{A}\times\mathcal{A}$ has a natural $\mathcal{M}$-category structure obtained by the identification
$\mathrm{Hom}\left(\left(A_{0},A_{1}\right),\left(B_{0},B_{1}\right)\right)=\mathrm{Hom}\left(A_{0},B_{0}\right)\times\mathrm{Hom}\left(A_{1},B_{1}\right)\text{.}$
Likewise, the category of arrows $\mathcal{A}^{\rightarrow}$ in $\mathcal{A}$ has a natural $\mathcal{M}$-category structure. More generally, for every finite category $J$, the category $\mathcal{A}^{J}$ of diagrams of shape $J$ in $\mathcal{A}$ is an $\mathcal{M}$-category.
Definition 2.8.
Let $\mathcal{M}$ be a quasi-abelian category of abelian groups.
  • An additive $\mathcal{M}$-category is an $\mathcal{M}$-category $\mathcal{A}$ that has a zero object and biproducts, and such that the biproduct functor $\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ is an $\mathcal{M}$-functor;
@@
  • A quasi-abelian $\mathcal{M}$-category is an additive $\mathcal{M}$-category that is also quasi-abelian, and such that the kernel and cokernel functors are $\mathcal{M}$-functors, which implies that all the finite limit and colimit functors are $\mathcal{M}$-functors;
  • A triangulated $\mathcal{M}$-category is an additive $\mathcal{M}$-category $\mathcal{T}$ that is also triangulated, and such that the shift functor $\Sigma$ on $\mathcal{T}$ is an $\mathcal{M}$-functor, and for al objects $A,B$ of $\mathcal{T}$, $\mathrm{Hom}\left(A,-\right)$ is a homological functor in the sense of [32, Definition 1.1.7] when regarded as a functor $\mathcal{T}\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$, and $\mathrm{Hom}\left(-,B\right)$ is a cohomological functor when regarded as a functor $\mathcal{T}^{\mathrm{op}}\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$.